Hamilton Company applies manufacturing overhead costs to products based on direct labor hours. The company estimates manufacturing overhead cost for the year to be $260,000 and direct labor hours to be 20,000. Actual overhead and actual direct labor hours for the year were $285,000 and 23,000 hours, respectively.

Required:

1. Compute over- or underapplied overhead.

2a. Which accounts will be affected by the over- or underapplied manufacturing overhead?

2b. Will the accounts be increased or decreased to adjust for the over- or underapplied manufacturing overhead?
Answer #1

1) Overhead rate = 260000/20000 = 13 per DLH

Applied overhead = 23000*13 = 299000

Actual overhead = 285000

Over applied overhead = 299000-285000 = 14000

2a) Cost of goods sold and manufacturing overhead.

2b) Manufacturing overhead account decreased